Maggie Mae's has approved a new soft drink to be distributed with its other flavored soft drinks. Management has estimated the additional fixed costs for its new soft drink line to be $140,000 and its standard wholesale price of $0.55 per unit will be applied. The associated labor and material costs per unit have been estimated to be $0.25.

How many units to break even? What are the break-even revenues?
Answer #1

When we consider BEP, we need the following formula:

BEP = FIXED COST / PRICE PER UNIT - VARIABLE COST

BEP = 140000 / (0.55 - 0.25)

BEP = 466666.67 OR 466667 UNITS

2. Break even revenue = breakeven quantity * price per unit

= 466667 * 0.55 = $256666.85